OpenStudy (anonymous):

find the missing terms of the arithmetic sequence 2, __, __, __, -0.4

OpenStudy (anonymous):

please help @whpalmer4

OpenStudy (whpalmer4):

Okay, which term numbers do we know? We know \(a_1\), and \(a_?\)

OpenStudy (anonymous):

ok is A1=2 and A=-0.4??

OpenStudy (anonymous):

sorry read it wrong it is A5

OpenStudy (whpalmer4):

what is the term number of -0.4? is it \(a_1, a_2, a_3, a_4,a_5,a_6,a_7,a_{9375234}\)?

OpenStudy (whpalmer4):

Right, it is \(a_5\). So what's the formula for the \(n\)th term of an arithmetic sequence?

OpenStudy (whpalmer4):

\[a_n = a_1 ...\]

OpenStudy (anonymous):

An=A1+(n-1)(d) right

OpenStudy (whpalmer4):

Yes. So we know \(a_1\) and \(a_5\). Can you use that information to find \(d\)?

OpenStudy (anonymous):

i think so what if we use A1=2 and n=5 and we leave d as d right

OpenStudy (whpalmer4):

go for it!

OpenStudy (anonymous):

well what i have is \[a _{n}=6(d)\] what do i do now

OpenStudy (whpalmer4):

no, you don't have that... you started with \[a_n= a_1 + (n-1)d\]then you plugged in n = 5 \[a_5 = a_1 + (5-1)d\]\[a_5 = a_1+4d\]\[a_5 = 2+4d\]\[-0.4 = 2+4d\] You just didn't realize that is what you did ;-)

OpenStudy (anonymous):

oh ok so what do we do next??

OpenStudy (whpalmer4):

Solve \[-0.4 = 2 + 4d\] for \(d\), silly :-)

OpenStudy (anonymous):

oh im retarded lol!!!! -_- the answer is d=-.6

OpenStudy (whpalmer4):

yes. I prefer to always put a 0 in front of the decimal point, just to make it a bit easier to see (and not forget). So what's the whole formula, and what are the missing terms?

OpenStudy (anonymous):

ok the formula is \[a_{n}=2+(n-1)(-0.6)\] and the missing terms are \[a_{2}, a_{3}, a_{4}\]

OpenStudy (whpalmer4):

right. what are the values of \(a_2, a_3, a_4\)?

OpenStudy (anonymous):

\[a_{2}=1.4, a_{3}=.8, a_{4}=.2\]

OpenStudy (whpalmer4):

very good.

OpenStudy (anonymous):

thank you so much!!!!

OpenStudy (whpalmer4):

the key here was to realize that even though we only knew two terms, we really knew 3 of the 4 things in the equation.

OpenStudy (whpalmer4):

Now, how about I bend your mind a bit with a problem of my own, which you might well encounter in your homework or on a test, but haven't asked me yet so far...

OpenStudy (whpalmer4):

You know two terms in an arithmetic sequence. One of them, \(a_4 = 9\). The other one, \(a_{10}=21\). What is the formula for the sequence?

OpenStudy (anonymous):

i have no clue!!!!! you just made my face go 0.o

OpenStudy (whpalmer4):

it's not as hard as it seems...

OpenStudy (anonymous):

i dont know how to start

OpenStudy (whpalmer4):

brb, nature calls

OpenStudy (anonymous):

lol ok

OpenStudy (whpalmer4):

okay, got the solution yet? :-)

OpenStudy (anonymous):

no i dont

OpenStudy (whpalmer4):

I'll show you two ways you could work it out, the "in your head" way, and the more formal way. which do you want first?

OpenStudy (anonymous):

um formal

OpenStudy (whpalmer4):

Okay. We know two terms. We don't know \(a_1\), and we don't know \(d\). \[a_4 = 9\]\[a_{10}=21\] Let's write the equations: \[a_4 = a_1+(n-1)d = a_1+(4-1)d = a_1+3d\]\[a_{10} = a_1+(n-1)d = a_1+(10-1)d = a_1+9d\] Think about that for a minute, and tell me if it gives you the answer...

OpenStudy (whpalmer4):

remember, we know the values of \(a_4\) and \(a_{10}\)...

OpenStudy (anonymous):

umm not sure will d=6

OpenStudy (whpalmer4):

ooh....you're on the right track...

OpenStudy (whpalmer4):

but going in the wrong direction :-)

OpenStudy (anonymous):

ok what do i do

OpenStudy (whpalmer4):

Let's write the equations with those values plugged in: \[21=a_1+6d\]\[\,\,\,\,9 = a_1+3d\]------------- What if we subtract those two equations from each other? Straight down the columns...

OpenStudy (anonymous):

12=6d

OpenStudy (whpalmer4):

wait, stop! I typed the wrong things... \[21 = a_1 + 9d\]\[\,\,\,\,9 = a_1+3d\] -----------------------------

OpenStudy (anonymous):

so then when we divide we will get d=2

OpenStudy (whpalmer4):

yes, you got the right answer, even though I typed the wrong equation :-)

OpenStudy (whpalmer4):

a case of two wrongs making a right :-)

OpenStudy (whpalmer4):

Next time someone tells you that two wrongs don't make a right, you can tell them you've seen at least one case where they did :-)

OpenStudy (anonymous):

i knew what you were trying to say so i used the original equations to find the answer

OpenStudy (whpalmer4):

Ah, good for you! So, the "in your head" way is to observe that \(a_{10}\) is 6 terms on from \(a_{4}\) in the sequence, so the difference between the two must be 6 * the common difference. 21 - 9 = 12, and 12/6 = 2, so the common difference is 2.

OpenStudy (anonymous):

i like that way better it is easier lol so hey will you help me with one more question if i make it a new question

OpenStudy (whpalmer4):

Kind of neat how that works, isn't it? At first, your reaction is "how could we possibly figure that out?" and then when you do, it's more like "how could we possibly not have realized that right away?" :-)

OpenStudy (whpalmer4):

I'm always happy to help you...